On September 30, World Co. borrowed $1,000,000 on a 9% note payable. World paid the first of four quarterly payments of $264,200 when due on December 30. In its December 31, balance sheet, what amount should World report as note payable? $735,800 $750,000 $825,800 $758,300
Answer #1

Interest on notes payable =1000000*9%*3/12 =22500

Out of payment of 264200, interest portion was 22500, the rest amount is principal 241700

Carrying value is thus =1000000-241700 = 758,300